# VisTabNet: Adapting Vision Transformers for Tabular Data

### v1

@@ Abstract
Although deep learning models have had great success in natural language processing and computer vision, we do not observe comparable improvements in the case of tabular data, which is still the most common data type used in biological, industrial and financial applications. In particular, it is challenging to transfer large-scale pre-trained models to downstream tasks defined on small tabular datasets. To address this, we propose VisTabNet – a cross-modal transfer learning method, which allows for adapting Vision Transformer (ViT) with pre-trained weights to process tabular data. By projecting tabular inputs to patch embeddings acceptable by ViT, we can directly apply a pre-trained Transformer Encoder to tabular inputs. This approach eliminates the conceptual cost of designing a suitable architecture for processing tabular data, while reducing the computational cost of training the model from scratch. Experimental results on multiple small tabular datasets (less than 1k samples) demonstrate VisTabNet’s superiority, outperforming both traditional ensemble methods and recent deep learning models. The proposed method goes beyond conventional transfer learning practice and shows that pre-trained image models can be transferred to solve tabular problems, extending the boundaries of transfer learning.
@@ 1 Introduction
Deep learning achieved tremendous success in various domains, including natural language processing (NLP) [33], computer vision (CV) [16], or reinforcement learning (RL) [20]. By feeding modern neural network architectures with a large amount of data, we can generate a model that not only solves the underlying task but also can be transferred to downstream tasks at low conceptual and computational cost. Transformers have recently become one of the most prominent neural architectures used in NLP [31] and CV [7]. By applying a self-attention mechanism, transformers allow capturing global dependencies across the entire text or image, demonstrating remarkable improvements over convolutional and recurrent neural networks. In consequence, it is not surprising that several works focus on applying transformers beyond CV and NLP domains.
@@
In real-world applications, one of the most common data types is tabular data, comprising samples (rows) with the same set of features (columns). Tabular data is used in many fields, including biology [29], medicine [25], finance [5], manufacturing [22], and many other applications that are based on relational databases. According to recent reports, data science and machine learning developers work with tabular data as often as with texts or images [^1] https://www.statista.com/statistics/1241924/worldwide-software-developer-data-uses/. This proportion is partially reflected in the Kaggle service [^2] Statistics gatherer in 2023 from https://www.kaggle.com/datasets, where 6 688 of available datasets are tagged as ”tabular”, 4 908 datasets contain the tag ”image” and 178 datasets are tagged as ”text”.
Although practical applications are dominated by tabular data, deep learning models do not confirm significant improvements over traditional ensemble models based on decision trees, such as XGBoost and Random Forests, in this domain [^3] https://www.kaggle.com/kaggle-survey-2021 [3, 15]. Due to its heterogeneous nature and intricate feature relationships, the complexity of tabular data poses challenges to existing deep learning models [11, 19]. Moreover, the small sample size of many tabular datasets is extremely problematic to handle by deep models. Inspired by the success of transformers in CV and NLP, analogical architectures have been crafted for tabular data [9], but it is not obvious how to pre-train and transfer these models to downstream tasks [34]. Moreover, creating and training transformer-based models for every single domain requires high conceptual and computational cost and may not work for small datasets, which are, however, ubiquitous for tabular data.
In this paper, we follow an emerging trend in deep learning and focus on transferring as much knowledge from other modalities as possible to construct a tabular transformer. Inspired by the idea of transfer learning, we reuse the Vision Transformer (ViT) to solve tasks defined for tabular data; see Figure 1. In contrast to the typical reasoning behind transferring a feature extractor inside the same domain, we explore cross-modal transfer. More precisely, we take the Transformer Encoder of ViT pre-trained on image data and introduce an adaptation network for mapping tabular inputs to the form acceptable by ViT Encoder. By aligning transformed tabular inputs with patch embeddings, we can use pre-trained ViT encoder to build abstract and meaningful representation of tabular data. In the resulting VisTabNet model, we avoid the conceptual cost of designing the correct architecture, as well as the computational cost of training the model from scratch. This, in turn, allows us to train a large and complex model relatively quickly and using a small amount of data.
VisTabNet has been validated using multiple conventional small tabular datasets. It obtained the best overall score among various recent deep learning methods and the well-established shallow ensembles (Table 1). We also show that VisTabNet transfers effectively using only a few examples per class (Figure 2), which further confirms that cross-modal transfer is more effective than training deep networks from scratch. Our evaluation is supported by the extensive experimental analysis of the main building blocks of VisTabNet (Section 4.2), which provides better insight into cross-modal transfer and the practical use of VisTabNet.
@@ 4.1 Tabular Data Classification
We consider small datasets retrieved from the UCI repository, which are summarized in Table 4. Small datasets are the most challenging case for deep learning methods, but thanks to applying transfer learning principle, VisTabNet is capable of reducing the overfitting issue.
@@ 4.2 Analysis of VisTabNet components
The heatmap presented in Figure 4 shows the average MCC scores obtained by transferring the part of the ViT encoder ranging from the ”start layer” to the ”end layer”. The standard variant of VisTabNet visualized in the top right corner (start = 0, end = 12) results in the MCC value of 0.75, which is 3 percentage points lower than the best score obtained going from the 5th to 12th layer. It can be observed that projecting data onto the 5th layer gives generally very good results. While initial layers of the ViT encoder are pre-trained to process linearly transformed image patches, later layers work on more abstract representations, which may better suit to tabular data. In consequence, if we start from the first layer of the ViT encoder, we need to apply many transformation (ending in the 12th layer) to get meaningful representation. Starting with more general transformations defined by the 5th layer, we do not need to apply so deep networks, which provide good balance between performance (measured by the MCC) and computation time (network reduction led to a 2-3 times reduction in computational time). Detailed results on individual datasets can be found in Figures 5
@@ Acknowledgements
This research has been supported by the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the Priority Research Area DigiWorld under the Strategic Programme Excellence Initiative at Jagiellonian University. This research was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169.

### v2

@@ Abstract
Although deep learning models have had great success in natural language processing and computer vision, we do not observe comparable improvements in the case of tabular data, which is still the most common data type used in biological, industrial and financial applications. In particular, it is challenging to transfer large-scale pre-trained models to downstream tasks defined on small tabular datasets. To address this, we propose VisTabNet – a cross-modal transfer learning method, which allows for adapting Vision Transformer (ViT) with pre-trained weights to process tabular data. By projecting tabular inputs to patch embeddings acceptable by ViT, we can directly apply a pre-trained Transformer Encoder to tabular inputs. This approach eliminates the conceptual cost of designing a suitable architecture for processing tabular data, while reducing the computational cost of training the model from scratch. Experimental results on multiple small tabular datasets (less than 1k samples) demonstrate VisTabNet’s superiority, outperforming both traditional ensemble methods and recent deep learning models. The proposed method goes beyond conventional transfer learning practice and shows that pre-trained image models can be transferred to solve tabular problems, extending the boundaries of transfer learning. We share our example implementation as a GitHub repository available at https://github.com/wwydmanski/VisTabNet.
@@ 1 Introduction
Deep learning has achieved tremendous success in various domains, including natural language processing (NLP) [34], computer vision (CV) [16], and reinforcement learning (RL) [20]. Transformers, in particular, have become one of the most prominent neural architectures in NLP [31] and CV [7], demonstrating remarkable improvements through their self-attention mechanism that captures global dependencies across inputs. In consequence, it is not surprising that several works focus on applying transformers beyond CV and NLP domains.
@@
In real-world applications, tabular data remains one of the most common data types, used extensively in biology [29], medicine [25], finance [5], and manufacturing [22]. Recent reports indicate that data science practitioners work with tabular data as frequently as with texts or images [^1] https://www.statista.com/statistics/1241924/worldwide-software-developer-data-uses/. This is reflected in Kaggle’s dataset distribution [^2] Statistics gatherer in 2023 from https://www.kaggle.com/datasets, where 6,688 datasets are tagged as ”tabular”, compared to 4,908 tagged as ”image” and 178 as ”text”.
Despite the prevalence of tabular data in practical applications, deep learning models have yet to demonstrate significant improvements over traditional ensemble methods like XGBoost and Random Forests in this domain [3, 15]. The heterogeneous nature of tabular data and typically small sample sizes pose particular challenges for deep learning approaches [11, 19]. While various deep learning approaches, including transformer architectures [9] and hypernetworks [32], have been adapted for tabular data, pre-training and transferring these models to downstream tasks remains challenging [35].
In this paper, we propose a novel approach to cross-modal transfer learning, reusing the Vision Transformer (ViT) for tabular data tasks; see Figure 1. In contrast to the typical reasoning behind transferring a feature extractor inside the same domain, we explore cross-modal transfer. More precisely, we take the Transformer Encoder of ViT pre-trained on image data and introduce an adaptation network that maps tabular inputs to a form compatible with the pre-trained ViT Encoder, enabling the construction of meaningful representations while avoiding the computational cost of training from scratch. Our VisTabNet model demonstrates superior performance across multiple conventional tabular datasets (Table 1) and shows effective few-shot transfer capabilities (Figure 2). Through extensive experimental analysis (Section 4.2), we provide insights into the effectiveness of cross-modal transfer and practical applications of VisTabNet.
@@ 4.1 Tabular Data Classification
We consider small datasets retrieved from the UCI repository, which are summarized in Table 1 in Appendix A [^3] Available at https://github.com/wwydmanski/VisTabNet/blob/main/Appendix.pdf. Small datasets are the most challenging case for deep learning methods, but thanks to applying transfer learning principle, VisTabNet is capable of reducing the overfitting issue.
@@ 4.2 Analysis of VisTabNet components
The heatmap presented in Figure 4 shows the average MCC scores obtained by transferring the part of the ViT encoder ranging from the ”start layer” to the ”end layer”. The standard variant of VisTabNet visualized in the top right corner (start = 0, end = 12) results in the MCC value of 0.75, which is 3 percentage points lower than the best score obtained going from the 5th to 12th layer. It can be observed that projecting data onto the 5th layer gives generally very good results. While initial layers of the ViT encoder are pre-trained to process linearly transformed image patches, later layers work on more abstract representations, which may better suit to tabular data. In consequence, if we start from the first layer of the ViT encoder, we need to apply many transformation (ending in the 12th layer) to get meaningful representation. Starting with more general transformations defined by the 5th layer, we do not need to apply so deep networks, which provide good balance between performance (measured by the MCC) and computation time (network reduction led to a 2-3 times reduction in computational time). Detailed results on individual datasets can be found in Figure 1 of Appendix B.
@@ Acknowledgements
This research has been supported by the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the Priority Research Area DigiWorld under the Strategic Programme Excellence Initiative at Jagiellonian University. This research was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016610. The work of Witold Wydmański is supported by the Ministry of Science grant no. PN/01/0195/2022. We would like to thank Łukasz Struski for invaluable discussions and his help at the initial stage of this work.